# Context is Key: A Benchmark for Forecasting with Essential Textual Information

### v3

@@ Abstract
Forecasting is a critical task in decision-making across numerous domains. While historical numerical data provide a start, they fail to convey the complete context for reliable and accurate predictions. Human forecasters frequently rely on additional information, such as background knowledge and constraints, which can efficiently be communicated through natural language. However, in spite of recent progress with LLM-based forecasters, their ability to effectively integrate this textual information remains an open question. To address this, we introduce “Context is Key” (CiK), a time-series forecasting benchmark that pairs numerical data with diverse types of carefully crafted textual context, requiring models to integrate both modalities; crucially, every task in CiK requires understanding textual context to be solved successfully. We evaluate a range of approaches, including statistical models, time series foundation models, and LLM-based forecasters, and propose a simple yet effective LLM prompting method that outperforms all other tested methods on our benchmark. Our experiments highlight the importance of incorporating contextual information, demonstrate surprising performance when using LLM-based forecasting models, and also reveal some of their critical shortcomings. This benchmark aims to advance multimodal forecasting by promoting models that are both accurate and accessible to decision-makers with varied technical expertise. The benchmark can be visualized at https://servicenow.github.io/context-is-key-forecasting/v0/.
@@ 1 Introduction
The prediction of future states of the world is a cornerstone of decision making (Hyndman & Athanasopoulos, 2018) and intelligence (Wang, 2019). Articulated as time-series forecasting, this problem pervades much of science and commerce.
@@ 3.1 Domains and Numerical Data Sources
Memorization mitigation:  Building tasks with publicly available data introduces contamination risk: pretrained LLMs and time-series foundation models may have memorized portions of the data, potentially inflating evaluation performance. We employ several mitigation strategies. First, we prioritize live data sources that are continuously updated, such as Chen et al. (2001) and Ville de Montréal (2020), ensuring the data is collected after the training cut-off dates of the models that we evaluate. Second, where applicable, we use derived series that are not directly available in the raw data, such as incident logs converted into time series (Ville de Montréal, 2020). Finally, as a last resort, we apply minor transformations, such as adding noise or shifting timestamps, but use these sparingly to avoid misalignment between common-sense knowledge (e.g., holiday dates) and the numerical data. The exact mitigation methods used, per data source, are given in Sec. A.1.
@@ 5.2 Methods
LLM-based Forecasters:  We consider two prompt-based approaches. We propose “ direct prompt ”, a simple approach where we instruct the model to directly output a forecast as a structured output for all of the required timestamps (see Sec. D.1 for more details). We also evaluate LLM Processes (llmp; Requeima et al. (2024)), a method which autoregressively prompts the LLM multiple times to output a forecast. For each of these, we evaluate a variety of LLMs with diverse architectures and sizes, such as GPT-4o , GPT-4o-mini (Achiam et al., 2023), Qwen-2.5-{0.5B, 1.5B, 7B}  (Yang et al., 2024), Mixtral-8x7B (Jiang et al., 2024), Llama-3-{8B, 70B}  (Dubey et al., 2024), Llama-3.1-405B (Dubey et al., 2024). [^1] For llmp, we do not consider Llama-3.1-405B and GPT models as llmp requires loading model weights into memory, which is infeasible due to resource limitations and confidentiality of the respective models (see Sec. D.2.2 for more details). We also evaluate multimodal forecasting models, UniTime  (Liu et al., 2024c) and Time-LLM (ETTh1)  (Jin et al., 2024) each trained according to their respective authors’ guidelines (details in Sec. D.3). For all of these approaches, inference is performed zero-shot on the benchmark and we compare their performance with and without the natural language context.
Quantitative Forecasting Models:  We also evaluate a number of models that process only numerical data, but cannot integrate natural language. We evaluate Exponential Smoothing  (Gardner Jr. (1985)), ETS  (Hyndman et al., 2008), and ARIMA  (Box et al., 2015), three simple, but time-tested statistical approaches. We also evaluate four state-of-the-art TSFMs: Lag-Llama  (Rasul et al., 2023), Chronos  (Ansari et al., 2024), Moirai  (Woo et al., 2024), and TimeGEN  (Garza et al., 2023). We fit Exponential Smoothing , ETS and ARIMA to the individual numerical history of each task instance, which is the same input that the TSFMs process to produce forecasts zero-shot.
@@ 5.3 Results on CiK
On the other hand, Figure 5 (right) shows that LLM-based forecasters, when evaluated without context, no longer dominate the quantitative forecasting models. However, some LLM-based forecasters remain surprisingly competitive. For instance, multiple Llama-3 (llmp) models outperform at least $5$ of the quantitative models on the majority of tasks. The extended results in Sec. C.1 further substantiate this. In contrast, models such as Llama-3.1-405B-Inst ( direct prompt ) and GPT-4o ( direct prompt ) show significantly weaker forecasting performance without context. This suggests that such models are especially preferable in cases where context is available. This is also reflected in their aggregate scores without context (in App. Tab. 2).
Comparing llmp and direct prompt :  Clear patterns emerge when comparing these strategies. Figure 5 (right) shows that without context, llmp models exhibit stronger numerical forecasting performance than direct prompt models. This advantage likely stems from llmp ’s closer alignment with the forecasting task: llmp simply prompts the LLM to autoregressively predict the next value in the time series, a task well suited to base models with no instruction tuning. In contrast, direct prompt requires output forecasts to be structured, which relatively complicates the task.
This line of reasoning leads us to another observation on the impact of instruction tuning; as reflected in Tab. 1 and Figure 5, instruction tuning appears to generally degrade llmp performance. Llama-3 models showing a twofold decrease in performance after tuning—a behavior previously observed by Gruver et al. (2024). Interestingly, instruction tuning does not degrade the performance of Mixtral-8x7B . Finally, whereas llmp mostly suffers from instruction tuning, direct prompt requires forecasts to be produced in a specific structure, a skill that instruction-tuned models might be better at (see Sec. D.1.1 for details).
@@
Figure 6: RCRPS with and without context (log scale, lower is better). Full bars show performance with context; striped bars show performance without. In general, larger models outperform smaller models and benefit much more from context. direct prompt models all improve with context, other than Qwen-2.5-0.5B-Instruct . For llmp, larger models benefit from context, but smaller models fail to do so and perform worse in general.
@@ 5.4 Error Analysis
Foundation models are known to make mistakes or return factually inaccurate information (Bommasani et al., 2021). We find that models occasionally return forecasts that miss the ground truth by a large margin. We use the term significant failure to denote a forecast that over or undershoots by at least $500\%$ of the range of the ground truth; we clip the RCRPS of such instances to 5 to avoid them disproportionately skewing the aggregate score. Despite this, such significant failures impact the results in Tab. 1: GPT-4o with direct prompt , while emerging as a top-performer in most tasks (as reflected in its average rank), has a significantly higher aggregate RCRPS than models ranked worse, such as Mixtral-8x7B with llmp. As an example, GPT-4o with direct prompt fails significantly in a task with a context involving scientific notation (see Figure 25 and more examples in Sec. C.12). Notably, while a model may generally achieve a high win rate, a few significant failures can dominate its aggregate performance, as observed for Mixtral-8x7B-Inst with direct prompt . We analyze this in detail in Sec. C.10. These findings underscore the need for future work to develop more robust models that can handle context effectively while avoiding significant failures.
@@ 5.5 Inference Cost
Figure 7: Comparison of average RCRPS (per Tab. 1), vs. the parameter count of each method(lower is better for both). The GPT family, as well as TimeGEN, are left out as there is no information on them about parameter count. The dashed line illustrates the Pareto front: models above and to the right of this line are dominated. Quantitative forecasters dominate the low-parameter regime, while LLM-based methods such as Qwen-2.5-7B-Inst ( direct prompt ) or Llama-3-70B (llmp) and 3.1-405B-Inst ( direct prompt ) offer superior performance for a higher parameter count.
A key practical aspect for forecasting applications is the inference time of models and their associated cost. Figure 7 shows that, while Llama-3.1-405B-Instruct performs the best on average, it comes at the cost of a significantly higher parameter count than the quantitative forecasters. This emphasizes that while LLMs can be powerful context-aware forecasters, they come with a steep computational cost, which highlights the need for efficient models that balance both accuracy and resource demands. Of note is also that many LLM-based forecasters are pareto-dominated quantitative forecasters such as Lag-Llama and Chronos . This suggests that, beyond the ability to simply process text, a careful choice of prompting strategy and LLM is crucial for Pareto efficiency. Further, due to their high parameter count, the LLMs have inference times that are orders of magnitude longer than quantitative forecasters which are far more efficient for sustained usage (see App. Figure 13 for a comparison). LLMs require significant computational power, making them unsuitable for real-world practical forecasting at scale where speed and cost matter. However, additional research could improve the efficiency of context-aided forecasting models to match the cost-effectiveness of traditional models.
@@ 6 Related Work
Repurposing LLMs for Forecasting  A natural approach to context-aided forecasting is to build methods based on LLMs. Xue & Salim (2023) showed that forecasting could be framed as a question-answering problem. Subsequently, Gruver et al. (2024) and Requeima et al. (2024) showed that LLMs could generate accurate forecasts with sequence completion, and that textual side-information could be used to influence forecasts. However, their analysis is limited to illustrative examples rather than a comprehensive evaluation. Other approaches have incorporated time series into pretrained LLMs (Jin et al., 2024; Liu et al., 2024c; Zhang et al., 2024) by introducing special tokens used to represent patched time series patterns, or modifying their encoders to account for time series data (Jia et al., 2024). While these methods show promising results, their evaluations primarily rely on datasets where the contextual information is not guaranteed to improve forecasts over numerical data alone. As a result, it remains unclear whether their performance is due by accurate numerical forecasting or by effectively incorporating context; this shortcoming motivates our investigation into this question.
@@ 7 Discussion
Limitations:  While our benchmark provides valuable insights into the integration of contextual information in time series forecasting, it is important to acknowledge its limitations. Our study limits itself to the natural language modality for context, and excludes multivariate time series scenarios. Although we deliberately designed the tasks to assess how well forecasting models can integrate contextual information, our benchmark does not evaluate whether models can leverage latent relationships that might elude human observation. While we have taken steps to mitigate memorization concerns, as discussed in Sec. 3.1, achieving absolute certainty in this regard is challenging without strictly held-out data.
Future work:  There are several promising avenues for future work. Extensions to CiK could include multivariate forecasting tasks, or tasks that incorporate other modalities such as images, databases or spatiotemporal data. Tasks that deliberately challenge context length limitations or probe specific weaknesses of language models would also be valuable additions. Methods to improve the automatic generation of large, high-quality datasets for context-aided forecasting could complement CiK. Furthermore, this benchmark strongly motivates research into developing more accurate and efficient multimodal forecasting models, which it is well-positioned to support. Lastly, as models become more robust, they could be integrated into agentic systems with conversational interfaces, allowing forecasts to be augmented with human expertise and automatically retrieved information. Such advancements would represent a significant step toward automating and democratizing access to powerful forecasting tools.
@@
This paper’s goal is to advance the development and refinement of forecasting methods which can leverage contextual information. It can thus increase the likelihood that such methods will be adopted by various organizations. Therefore, this paper shares the positive and negative impacts that such methods will have. Such methods could increase forecasting accuracy and efficiency, increasing the ability to foresee future courses and plan in consequence. A secondary effect is an increased democratization of access to high-quality forecasts. On the flip side, this democratization could increase reliance on automated methods for decision-making purposes.
@@ C.2 Full results partitioned by types of context
    Llama-3.1-405B-Inst 0.159 $\pm$ 0.008 7.455 $\pm$ 0.585 0.174 $\pm$ 0.010 0.146 $\pm$ 0.001 0.066 $\pm$ 0.004 0.168 $\pm$ 0.010 0.398 $\pm$ 0.045
    Llama-3-70B-Inst 0.286 $\pm$ 0.004 14.082 $\pm$ 0.194 0.336 $\pm$ 0.006 0.180 $\pm$ 0.003 0.185 $\pm$ 0.006 0.230 $\pm$ 0.005 0.629 $\pm$ 0.019
    Llama-3-8B-Inst 0.461 $\pm$ 0.008 26.903 $\pm$ 0.558 0.571 $\pm$ 0.011 0.313 $\pm$ 0.008 0.241 $\pm$ 0.015 0.263 $\pm$ 0.003 0.531 $\pm$ 0.005
    Mixtral-8x7B-Inst 0.523 $\pm$ 0.023 31.390 $\pm$ 0.393 0.724 $\pm$ 0.037 0.236 $\pm$ 0.002 0.185 $\pm$ 0.006 0.384 $\pm$ 0.028 0.875 $\pm$ 0.128
    Qwen-2.5-7B-Inst 0.289 $\pm$ 0.003 21.792 $\pm$ 0.725 0.289 $\pm$ 0.004 0.176 $\pm$ 0.003 0.281 $\pm$ 0.007 0.238 $\pm$ 0.002 0.525 $\pm$ 0.003
    Qwen-2.5-1.5B-Inst 0.575 $\pm$ 0.014 30.634 $\pm$ 0.834 0.682 $\pm$ 0.023 0.284 $\pm$ 0.006 0.358 $\pm$ 0.009 0.453 $\pm$ 0.004 1.270 $\pm$ 0.009
    Qwen-2.5-0.5B-Inst 0.464 $\pm$ 0.012 26.254 $\pm$ 0.514 0.610 $\pm$ 0.019 0.165 $\pm$ 0.004 0.221 $\pm$ 0.011 0.483 $\pm$ 0.016 0.429 $\pm$ 0.006
    GPT-4o 0.276 $\pm$ 0.010 8.609 $\pm$ 0.402 0.220 $\pm$ 0.007 0.118 $\pm$ 0.001 0.113 $\pm$ 0.001 0.258 $\pm$ 0.012 0.858 $\pm$ 0.053
    GPT-4o-mini 0.353 $\pm$ 0.022 16.463 $\pm$ 0.463 0.474 $\pm$ 0.035 0.139 $\pm$ 0.002 0.139 $\pm$ 0.002 0.343 $\pm$ 0.029 0.644 $\pm$ 0.128

### v4

@@ Abstract
Forecasting is a critical task in decision-making across numerous domains. While historical numerical data provide a start, they fail to convey the complete context for reliable and accurate predictions. Human forecasters frequently rely on additional information, such as background knowledge and constraints, which can efficiently be communicated through natural language. However, in spite of recent progress with LLM-based forecasters, their ability to effectively integrate this textual information remains an open question. To address this, we introduce “Context is Key” (CiK), a time series forecasting benchmark that pairs numerical data with diverse types of carefully crafted textual context, requiring models to integrate both modalities; crucially, every task in CiK requires understanding textual context to be solved successfully. We evaluate a range of approaches, including statistical models, time series foundation models, and LLM-based forecasters, and propose a simple yet effective LLM prompting method that outperforms all other tested methods on our benchmark. Our experiments highlight the importance of incorporating contextual information, demonstrate surprising performance when using LLM-based forecasting models, and also reveal some of their critical shortcomings. This benchmark aims to advance multimodal forecasting by promoting models that are both accurate and accessible to decision-makers with varied technical expertise. The benchmark can be visualized at https://servicenow.github.io/context-is-key-forecasting/v0/.
@@ 1 Introduction
The prediction of future states of the world is a cornerstone of decision making (Hyndman & Athanasopoulos, 2018) and intelligence (Wang, 2019). Articulated as time series forecasting, this problem pervades much of science and commerce.
@@ 3.1 Domains and Numerical Data Sources
Memorization mitigation:  Building tasks with publicly available data introduces contamination risk: pretrained LLMs and time series foundation models may have memorized portions of the data, potentially inflating evaluation performance. We employ several mitigation strategies. First, we prioritize live data sources that are continuously updated, such as Chen et al. (2001) and Ville de Montréal (2020), ensuring the data is collected after the training cut-off dates of the models that we evaluate. Second, where applicable, we use derived series that are not directly available in the raw data, such as incident logs converted into time series (Ville de Montréal, 2020). Finally, as a last resort, we apply minor transformations, such as adding noise or shifting timestamps, but use these sparingly to limit their potential impact on the tasks, and to avoid misalignment between common-sense knowledge (e.g., holiday dates) and the numerical data. The exact mitigation methods used, per data source, together with the number of tasks on which they are applied are given in Sec. A.1.
@@ 5.2 Methods
LLM-based Forecasters:  We consider two prompt-based approaches. We propose “ direct prompt ”, a simple approach where we instruct the model to directly output a forecast as a structured output for all of the required timestamps (see Sec. D.1 for more details). We also evaluate LLM Processes (llmp; Requeima et al. (2024)), a method which autoregressively prompts the LLM multiple times to output a forecast. For each of these, we evaluate a variety of LLMs with diverse architectures and sizes, such as GPT-4o , GPT-4o-mini (Achiam et al., 2023), Qwen-2.5-{0.5B, 1.5B, 7B}  (Yang et al., 2024), Mixtral-8x7B (Jiang et al., 2024), Llama-3-{8B, 70B}  (Dubey et al., 2024), Llama-3.1-405B (Dubey et al., 2024). [^1] For llmp, we do not consider Llama-3.1-405B and GPT models as llmp requires loading model weights into memory, which is infeasible due to resource limitations and confidentiality of the respective models (see Sec. D.2.2 for more details). We also evaluate multimodal forecasting models, ChatTime (Wang et al., 2025) (zero-shot), UniTime  (Liu et al., 2024c) and Time-LLM (ETTh1)  (Jin et al., 2024) (trained according to their respective authors’ guidelines) (details in Sec. D.4). For all of these approaches, inference is performed zero-shot on the benchmark and we compare their performance with and without the natural language context.
Quantitative Forecasting Models:  We also evaluate a number of models that process only numerical data, but cannot integrate natural language. We evaluate Exponential Smoothing  (Gardner Jr. (1985)), ETS  (Hyndman et al., 2008), and ARIMA  (Box et al., 2015), three simple, but time-tested statistical approaches. We also evaluate four state-of-the-art TSFMs: Lag-Llama  (Rasul et al., 2023), Chronos  (Ansari et al., 2024), Moirai  (Woo et al., 2024) and TimeGEN  (Garza et al., 2023). We fit Exponential Smoothing , ETS and ARIMA to the individual numerical history of each task instance, which is the same input that the TSFMs process to produce forecasts zero-shot.
@@ 5.3 Results on CiK
On the other hand, Figure 5 (right) shows that LLM-based forecasters, when evaluated without context, no longer dominate the quantitative forecasting models. However, some LLM-based forecasters remain surprisingly competitive. For instance, multiple Llama-3 (llmp) models outperform at least $5$ of the quantitative models on the majority of tasks. The extended results in Sec. C.1 further substantiate this. In contrast, models such as Llama-3.1-405B-Inst ( direct prompt ) and GPT-4o ( direct prompt ) show significantly weaker forecasting performance without context. This suggests that such models are especially preferable in cases where context is available. This is also reflected in their aggregate scores without context (in Tab. 3).
Comparing llmp and direct prompt :  Figure 5 (right) shows that without context, llmp models exhibit stronger numerical forecasting performance than direct prompt models. This advantage likely stems from llmp ’s closer alignment with the forecasting task: llmp simply prompts the LLM to autoregressively predict the next value in the time series, a task well suited to base models with no instruction tuning. In contrast, direct prompt requires output forecasts to be structured, which relatively complicates the task.
This line of reasoning leads us to another observation on the impact of instruction tuning; as reflected in Tab. 1 and Figure 5, instruction tuning appears to generally degrade llmp performance. Llama-3 models show a twofold decrease in performance after instruction tuning, a behavior previously observed by Gruver et al. (2024). Interestingly, instruction tuning does not degrade the performance of Mixtral-8x7B . Finally, whereas llmp mostly suffers from instruction tuning, direct prompt requires forecasts to be produced in a specific structure, a skill that instruction-tuned models might be better at (see Sec. D.1.1 for details).
@@
Figure 6: RCRPS with and without context (log scale, lower is better). Full bars show performance with context; striped bars show performance without context. In general, larger models outperform smaller models and benefit much more from context. direct prompt models all improve with context, other than Qwen-2.5-0.5B-Instruct . For llmp, larger models benefit from context, but smaller models fail to do so and perform worse in general.
@@ 5.4 Error Analysis
Foundation models are known to make mistakes or return factually inaccurate information (Bommasani et al., 2021). We find that models occasionally return forecasts that miss the ground truth by a large margin. We use the term significant failure to denote a forecast that over or undershoots by at least $500\%$ of the range of the ground truth; we clip the RCRPS of such instances to 5 to avoid them disproportionately skewing the aggregate score. Despite this, such significant failures impact the results in Tab. 1: GPT-4o with direct prompt , while emerging as a top-performer in most tasks (as reflected in its average rank), has a significantly higher aggregate RCRPS than models ranked worse, such as Mixtral-8x7B with llmp. As an example, GPT-4o with direct prompt fails significantly in a task with a context involving scientific notation (see Figure 25 and more examples in Sec. C.12). Notably, while a model may generally achieve a high win rate, a few significant failures can dominate its aggregate performance, as observed for GPT-4o with direct prompt . In the case of Qwen-2.5-0.5B-Inst with direct prompt , this leads to an aggregate RCRPS that is worse with context than without. We analyze this in detail in Sec. C.10. These findings underscore the need for future work to develop more robust models that can handle context effectively while avoiding significant failures.
@@ 5.5 Inference Cost
Figure 7: Comparison of average RCRPS (per Tab. 1), vs. the parameter count of each method(lower is better for both). The GPT family, as well as TimeGEN, are left out as there is no information on them about parameter count. The dashed line illustrates the Pareto front: models above and to the right of this line are dominated. Quantitative forecasters dominate the low-parameter regime, while LLM-based methods such as Qwen-2.5-7B-Inst ( direct prompt ), Llama-3-70B (llmp) and Llama-3.1-405B-Inst ( direct prompt ) offer superior performance for a higher parameter count.
A key practical aspect for forecasting applications is the inference time of models and their associated cost. Figure 7 shows that, while Llama-3.1-405B-Instruct performs the best on average, it comes at the cost of a significantly higher parameter count than the quantitative forecasters. This emphasizes that while LLMs can be powerful context-aware forecasters, they come with steep computational costs, which highlights the need for efficient models that balance both accuracy and resource demands. Of note is also that many LLM-based forecasters are pareto-dominated by quantitative forecasters such as Lag-Llama and Chronos . This suggests that, beyond the ability to process text, a careful choice of prompting strategy and LLM is crucial for Pareto efficiency. Further, due to their high parameter count, the LLMs have inference times that are orders of magnitude longer than quantitative forecasters, which are far more efficient for sustained usage (see Figure 13 for a comparison). LLMs require significant computational power, making them unsuitable for real-world practical forecasting at scale where speed and cost matter. However, additional research could improve the cost-effectiveness of context-aided forecasting models to match that of traditional models.
@@ 6 Related Work
Repurposing LLMs for Forecasting  A natural approach to context-aided forecasting is to build methods based on LLMs. Xue & Salim (2023) showed that forecasting could be framed as a question-answering problem. Subsequently, Gruver et al. (2024) and Requeima et al. (2024) showed that LLMs could generate accurate forecasts with sequence completion, and that textual side-information could be used to influence forecasts. However, their analysis is limited to illustrative examples rather than a comprehensive evaluation. Some works have explored the ability of LLMs to reason about time series (Chow et al., 2024; Kong et al., 2025; Aksu et al., 2024; Potosnak et al., 2024; Ye et al., 2024). Other approaches have incorporated time series into pretrained LLMs (Jin et al., 2024; Liu et al., 2024c; Zhang et al., 2024) by introducing special tokens used to represent patched time series patterns, or modifying their encoders to account for time series data (Jia et al., 2024). While these methods show promising results, their evaluations primarily rely on datasets where the contextual information is not guaranteed to improve forecasts over numerical data alone. As a result, it remains unclear whether their performance is due by accurate numerical forecasting or by effectively incorporating context; this shortcoming motivates our investigation into this question.
@@ 7 Discussion
Limitations:  While our benchmark provides valuable insights into the integration of contextual information in time series forecasting, it is important to acknowledge its limitations. Our study limits itself to the natural language modality for context, and excludes multivariate time series scenarios. Although we deliberately designed the tasks to assess how well forecasting models can integrate contextual information, our benchmark does not evaluate whether models can leverage latent relationships that might elude human observation. While we have taken steps to mitigate memorization concerns, as discussed in Sec. 3.1, achieving absolute certainty in this regard is challenging without strictly held-out data. Finally, the performance of dataset-specific methods such as ChatTime (Wang et al., 2024a), UniTime (Liu et al., 2024c) and Time-LLM (Jin et al., 2024) may improve in the presence of a dataset-specific training set.
Future work:  There are several promising avenues for future work. Extensions to CiK could include multivariate forecasting tasks, or tasks that incorporate other modalities such as images, databases or spatiotemporal data. Tasks that deliberately challenge context length limitations, probe specific weaknesses of language models or include domain-specific expert knowledge would also be valuable additions. More research is also needed to better understand what drives the catastrophic failures that LLMs exhibit. The analysis of catastrophic failures could benefit from searching for systematic failure patterns, or failures associated with specific linguistic patterns. Training datasets for context-aided forecasting would enable a better evaluation of dataset-specific methods. In fact, methods to improve the automatic generation of large, high-quality datasets for context-aided forecasting could complement CiK. Furthermore, this motivates research into developing more accurate and efficient multimodal forecasting models, which our benchmark is well-positioned to support. Other avenues of research include exploring different input/output structures for forecasting with LLMs, finetuning specialized LLMs for context-aided forecasting, allowing models to scale test-time computation, and compressing the context to reduce the required amount of computation. Lastly, as models become more robust, they could be integrated into agentic systems with conversational interfaces, allowing forecasts to be augmented with human expertise and automatically retrieved information. Such advancements would represent a significant step toward automating and democratizing access to powerful forecasting tools.
@@
This paper’s goal is to advance the development of forecasting methods which can leverage contextual information, which can increase the likelihood that such methods will be adopted by various organizations. Therefore, this paper shares the positive and negative impacts that such methods will have. Such methods could increase forecasting accuracy and efficiency, increasing the ability to foresee future courses and plan in consequence. A secondary effect is an increased democratization of access to high-quality forecasts. On the flip side, this democratization could increase reliance on automated methods for decision-making purposes.
@@ A.1 Data Sources
Table 2: Summary of Transformations Applied to Tasks Per Domain
Domain Number of tasks Transformations
Date shift Gaussian noise
Public Safety 26 None None
Traffic 11 None None
Mechanics 3 None None
Economics 3 None None
Synthetic 3 None None
Total 46 None None
Climatology 12 9 3
Energy 7 None 7
Retail 6 None 6
Total 25 9 16
@@
A summary of the number of tasks with either types of memorization strategy (shifting the dates by one day, or adding gaussian noise) is presented in Tab. 2.
@@ C.2 Full results partitioned by types of context
    Llama-3.1-405B-Inst 0.159 $\pm$ 0.008 7.967 $\pm$ 0.587 0.174 $\pm$ 0.010 0.146 $\pm$ 0.001 0.075 $\pm$ 0.005 0.164 $\pm$ 0.010 0.398 $\pm$ 0.045
    Llama-3-70B-Inst 0.286 $\pm$ 0.004 14.806 $\pm$ 0.201 0.336 $\pm$ 0.006 0.180 $\pm$ 0.003 0.194 $\pm$ 0.006 0.228 $\pm$ 0.004 0.629 $\pm$ 0.019
    Llama-3-8B-Inst 0.461 $\pm$ 0.008 28.013 $\pm$ 0.588 0.572 $\pm$ 0.011 0.313 $\pm$ 0.008 0.253 $\pm$ 0.017 0.262 $\pm$ 0.003 0.531 $\pm$ 0.005
    Mixtral-8x7B-Inst 0.523 $\pm$ 0.023 33.068 $\pm$ 0.415 0.723 $\pm$ 0.037 0.236 $\pm$ 0.002 0.241 $\pm$ 0.001 0.359 $\pm$ 0.028 0.875 $\pm$ 0.128
    Qwen-2.5-7B-Inst 0.290 $\pm$ 0.003 22.842 $\pm$ 0.733 0.290 $\pm$ 0.004 0.176 $\pm$ 0.003 0.287 $\pm$ 0.007 0.240 $\pm$ 0.002 0.525 $\pm$ 0.003
    Qwen-2.5-1.5B-Inst 0.575 $\pm$ 0.014 32.642 $\pm$ 0.864 0.684 $\pm$ 0.023 0.284 $\pm$ 0.006 0.370 $\pm$ 0.010 0.450 $\pm$ 0.004 1.270 $\pm$ 0.009
    Qwen-2.5-0.5B-Inst 0.463 $\pm$ 0.012 27.218 $\pm$ 0.530 0.609 $\pm$ 0.019 0.165 $\pm$ 0.004 0.218 $\pm$ 0.012 0.476 $\pm$ 0.015 0.429 $\pm$ 0.006
    GPT-4o 0.274 $\pm$ 0.010 8.631 $\pm$ 0.443 0.218 $\pm$ 0.007 0.118 $\pm$ 0.001 0.121 $\pm$ 0.001 0.250 $\pm$ 0.011 0.858 $\pm$ 0.053
    GPT-4o-mini 0.354 $\pm$ 0.022 17.565 $\pm$ 0.506 0.475 $\pm$ 0.035 0.139 $\pm$ 0.002 0.143 $\pm$ 0.002 0.341 $\pm$ 0.028 0.644 $\pm$ 0.128